I am having trouble getting the correct answer to this problem. The problem is
1/x+3 = -2x/x+5
This is how I worked it. The answer is -5/2, -1 but I keep getting -5/6
x+3(-2x) = x+5(1)
x-6x = x+5
-6x = 5
divide through by -6 and get -5/6.
What am I missing?

Found 2 solutions by unlockmath, mananth:
Answer by unlockmath(1688)

Hello,
You did good until the 3rd step.
We'll first get rid of the denominator so multiply each side by (x+3)(x+5) to get:
1(x+5) = -2x(x+3)
Expand it out to:
x+5=-2x^2-6x
Add 2x^2 and 6x to both sides to get:
2x^2+7x+5=0
This can be factored to:
(2x+5)(x+1)=0
Solve for x to get:
x=-1
x=-5/2
Make sense?
RJ
www.math-unlock.com

Answer by mananth(16946)

1/x+3 = -2x/x+5
..
x+5=-2x(x+3)
x+5=-2x^2-6x
-5
x+5-5 = -2x^2-6x-5
x=-2x^2-6x -5
2x^2+7x+5=0
2x^2+2x+5x+5=0
2x(x+1)+5(x+1)=0
(2x+5)(x+1)=0
x= -5/2 OR x=-1
...
m.ananth@hotmail.ca